OpenStudy (anonymous):

Rewrite the rational exponent as a radical. 5 to the 3 over 4 power, to the 2 over 3 power Answer Choices: A) the cube root of 5 squared B) the twelfth root of 5 C) the square root of 5 D)the cube root of 5 the fourth power I think it's A but im not sure, someone please help?

Nnesha (nnesha):

\[\huge\rm (5^\frac{ 5 }{ 4 })^\frac{ 2 }{ 3 }\] like this ?

OpenStudy (anonymous):

@Nnesha Yes.

Nnesha (nnesha):

well can you take a screenshot i think i typed it wrong

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@Nnesha did you get it?

Nnesha (nnesha):

weird question .... why do you think it's A ?

OpenStudy (anonymous):

@Nnesha When I worked out the problem I some how ended up with A

Nnesha (nnesha):

well i didn't get A i got*NONE* ORR i'm not paying attention

Nnesha (nnesha):

@freckles ???

Nnesha (nnesha):

you have to apply exponent rule \[\huge\rm (x^m)^n = x^{m \times n}\] multiply exponents and then reduce the fraction

OpenStudy (anonymous):

@Nnesha I did so, I did it 4 more times and got different answers each time. I have no idea anymore.

OpenStudy (freckles):

hey

OpenStudy (freckles):

hey @Nnesha he said to ...to the 3/4 power... but then he said yes when you put 5/4 instead...

OpenStudy (freckles):

\[(5^{\frac{3}{4})\frac{2}{3}} =5^{\frac{3}{4} \cdot \frac{2}{3}}\] first look at the fraction multiplication action going on there: \[\frac{3}{4} \cdot \frac{2}{3}\] can you simplify that?

OpenStudy (freckles):

basically you need to simplify 3/3 and 2/4

Nnesha (nnesha):

*facepalm* it's 3/4 >.<

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha \[\huge\rm (5^\frac{ 5 }{ 4 })^\frac{ 2 }{ 3 }\] like this ? \(\color{blue}{\text{End of Quote}}\) \(\color{blue}{\text{Originally Posted by}}\) @06292014 @Nnesha Yes. \(\color{blue}{\text{End of Quote}}\) ......

OpenStudy (anonymous):

wow my mistake guys :/ im so sorry

Nnesha (nnesha):

my bad.. thanks freckles

OpenStudy (freckles):

does @06292014 still need help? if so have you done the multiplication I asked about?

OpenStudy (anonymous):

@freckles I got 1/2

OpenStudy (freckles):

cool so you have 5^(1/2)

OpenStudy (freckles):

which is the same as square root of 5

OpenStudy (anonymous):

@freckles @Nnesha thank you so much both of you for your help!

OpenStudy (anonymous):

@Nnesha @freckles

OpenStudy (freckles):

look at the exponents do top exponent - bottom exponent

Nnesha (nnesha):

i can't zoomin in this pdf http://assets.openstudy.com/updates/attachments/557890d9e4b0636b8cc2974f-06292014-1433965330041-cuserspattydocumentsscreenshot.pdf and people like me barely can see this^ ;P

Nnesha (nnesha):

and write these down 'n ur notes \[\huge\rm x^m \times x^n = x^{m+n}\]\[\large\rm (x^m)^n =x^{m \times n}\]\[\large\rm \frac{ x^m }{ x^n }= x^{m-n}\]